OpenStudy (anonymous):

simplify the following problem using the distributive property : 2x+3y+2(x-y)-3x

OpenStudy (isaiah.feynman):

Simply distribute the 2 into x-y then add like terms.

OpenStudy (anonymous):

I got that much :)

OpenStudy (anonymous):

Please cont....

OpenStudy (isaiah.feynman):

Okay so you now have 2x+3y+2x-2y-3x. Right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So then I put the x's together 2x+2x-3x 4x-3x=1x Right?

OpenStudy (isaiah.feynman):

Correct! Except in maths we write 1x as just x. Now put the y's together.

OpenStudy (anonymous):

3y+2y=5y

OpenStudy (anonymous):

5y-3y=2y +x+2y

OpenStudy (anonymous):

So simplified is x + 2y

OpenStudy (anonymous):

?

OpenStudy (isaiah.feynman):

Actually its 3y-2y, which is y. This becomes x+y.

OpenStudy (anonymous):

so where did i get 5y from ?

OpenStudy (anonymous):

oh i see it is just x+y

OpenStudy (anonymous):

it is not 5y because you add it is 1y cause you subtract making the answer x+y !

OpenStudy (isaiah.feynman):

Yes. Correct.

OpenStudy (anonymous):

Yay Me ! Yay you Thanks So Much :)